# Measurement of zero-frequency fluctuations generated by coupling between Alfvén modes in the JET tokamak

### v1

@@ Abstract
Figure 4 shows the modulus of the bicoherence and the power spectrum at medium $P_{\text{ICRH}}$, calculated for a total of $63$ half-overlapping windows of period $T=0.5$ ms. The DBS beam reaches the cutoff surface at $\rho=0.355$ ($q\approx 1$), corresponding to the scattered wavenumber $k_{\perp}\rho_{s}=2.07$. The power spectrum $P_{s}(f)$ is broader than in the low-$P_{\text{ICRH}}$ phase, featuring fluctuations in the range $f\lesssim 800$ kHz, as well as spectral peaks at the local Alfvén gap frequency $f_{A}\approx 290$ kHz and a harmonic at $2f_{A}$. The spectral peaks appear for positive and negative frequencies, suggesting the presence of bidirectional Alfvénic activity. Careful analysis of the DBS microwave beam, pressure-constrained EFIT, and the $q=1$ constraint from ECE shows that the modes observed in figure 4 are well localized around $q=1$. This is consistent with previous observations of core-localized TAEs in TFTR [78, 79, 80] and tornado modes in JT-60U [49]. Such modes were predicted to exist at low shear [81], consistent with the low shear value in this JET experiment ($\hat{s}\lesssim 0.1$). The bicoherence in figure 4 displays broadband phase-matched relationships for multiples of the local Alfvén gap frequency $f_{A}$, the strongest of which are $(f_{1},f_{2})=\{(\pm f_{A},\pm f_{A})\}$ (amplitude $\approx 0.6$), which correspond to co-propagating TAEs that couple to generate a perturbation of frequency ±$\pm 2f_{A}$, as well as $(f_{1},f_{2})=\{(2f_{A},-f_{A}),(-f_{A},2f_{A})\}$, which correspond to perturbations of frequency $2f_{A}$ coupling to a counter-propagating perturbation at frequency $f_{A}$ that generates a co-propagating perturbation at $f_{A}$. The bicoherence also clearly exhibits phase-matched relationships between perturbations at multiples of $f_{A}$ and a zero-frequency fluctuation. To our knowledge, this is the first experimental demonstration in a magnetically confined plasma that Alfvén modes living inside the fundamental toroidicity-induced frequency gap (and higher-frequency gaps) beat with counter-propagating perturbations to generate a zero-frequency fluctuation. The zero-frequency fluctuations observed here can naturally be interpreted as zonal modes driven by Alfvén eigenmodes, which have been predicted analytically [16, 17, 18, 19] and numerically [20, 22, 23, 24, 12, 11, 14]. Such zonal modes can suppress ITG-driven turbulence, which would explain the increase of the ion temperature and its gradient (figure 2).
Figure 4: The same as figure 3, but at medium $P_{\text{ICRH}}$. (Left) The bicoherence displays phase-matched relationships for multiples of $f_{A}$ (vertical dashed lines). The strongest interactions are between frequencies ±$\pm f_{A}$ and ±$\pm 2f_{A}$. A weaker phase-matched relationship is observed between fluctuations at multiples of $f_{A}$ and a zero-frequency fluctuation. (Right) The power spectrum $P_{s}(f)$ displays spectral peaks at ±$\pm f_{A}$ and ±$\pm 2f_{A}$.
@@
Figure 5 shows the modulus of the bicoherence and the power spectrum at high $P_{\text{ICRH}}$, calculated for a total of $127$ half-overlapping windows of period $T=1$ ms. The DBS beam reaches the cutoff surface at $\rho=0.251$ ($q\approx 0.92$) for $k_{\perp}\rho_{s}=3.78$. The power spectrum is broadband for fluctuations in the range $f\lesssim 600$ kHz, but also displays spectral peaks at the local Alfvén gap frequency $f_{A}\approx 290$ kHz and its multiples. Similarly to figure 4, the bicoherence displays broadband phase-matched relationships for multiples of $f_{A}$. Importantly, the highest-amplitude phase matching is now for the frequency pairs $(f_{1},f_{2})=\{(2f_{A},0),(2f_{A},-2f_{A}),(3f_{A},0),(3f_{A},-3f_{A})\}$ (amplitude $\approx 0.5$), i.e., the pairs that involve a zero-frequency fluctuation. This is different from the medium-$P_{\text{ICRH}}$ phase, in which the phase-matched relationships involving the zero-frequency fluctuation were subdominant to other mode-mode interactions. Interestingly, the modes at the fundamental Alfvén gap frequency $f_{A}$ display a weaker phase-matched relationship with the zero-frequency fluctuation. Higher multiples of the Alfvén frequency show weaker phase-matched relationships, resulting in a peculiar grid structure.
@@
Figure 5: (Left) The bicoherence spectrum at high-$P_{\text{ICRH}}$ displays phase-matched relationships for multiples of the local Alfvén gap frequency $f_{A}$. The strongest phase matching is for frequency pairs involving a zero-frequency fluctuation: $(f_{1},f_{2})=\{(\pm 2f_{A},0),(\pm 3f_{A},0)\}$. (Centre) Zoom into the left pannel. (Right) The power spectrum $P_{s}(f)$ displays spectral peaks at $f_{A}$ and its multiples.
@@
Figures 4 and 5 confirm that Alfvén modes living inside the toroidicity-induced Alfvén-frequency gap (and higher-frequency gaps) beat with counter-propagating Alfvén modes to generate a strong zero-frequency fluctuation that is consistent with a zonal mode. When this behavior is observed, the total energy-confinement factor increases ($H_{89,P}>1$), as do the main-ion temperature and its gradient in the deep core, despite negligible direct ion heating. This shows the strong beneficial effect that Alfvén eigenmodes can have on the confinement of the thermal plasma. This is to be constrasted with the stable-Alfvén-mode case (low $P_{\text{ICRH}}$), which exhibited low ion temperature and temperature gradient, and $H_{89,P}<1$. This is the first experimental confirmation in a magnetically confined plasma of a zero-frequency fluctuation that is pumped by Alfvén modes, and is correlated with hotter-than-expected main ions. It suggests that zonal modes driven by Alfvén modes can suppress the turbulence and lead to an overall improvement of energy confinement in the plasma.

### v2

@@ Abstract
Figure 3 shows the power spectrum $P_{s}(f)$ (in arbitrary units [a.u.]) at low, medium, and high $P_{\text{ICRH}}$. At low $P_{\text{ICRH}}$, the DBS beam reaches the cutoff surface at $\rho=0.247$ (where $q\approx 1$), corresponding to the scattered wavenumber $k_{\perp}\rho_{s}=2.58$. The power spectrum $P_{s}(f)$ is broadband for fluctuations in the range $f\lesssim 200$ kHz characteristic of drift-wave turbulence [89, 90], and particularly, turbulence driven by the ion-temperature gradient mode (ITG) (note that non-broadband, drift-wave-like turbulence associated with other linear micro-instabilities is reported in the literature [91, 92], but appendix A shows that those are not relevant here). A peak at zero frequency is observed, consistent with many previous DBS measurements. At medium $P_{\text{ICRH}}$, the DBS beam reaches the cutoff surface at $\rho=0.355$ ($q\approx 1$), corresponding to the scattered wavenumber $k_{\perp}\rho_{s}=2.07$. The power spectrum $P_{s}(f)$ is broader than in the low-$P_{\text{ICRH}}$ phase, featuring fluctuations in the range $f\lesssim 800$ kHz, as well as spectral peaks at the local Alfvén gap frequency $f_{\text{TAE}}\approx 290$ kHz and a harmonic at $2f_{\text{TAE}}$. The spectral peaks appear for positive and negative frequencies, which could suggest the presence of bidirectional Alfvénic activity. Careful analysis of the DBS microwave beam, pressure-constrained EFIT, and the $q=1$ constraint from ECE shows that the modes observed in figure 3 are well localized around $q=1$. This is consistent with previous observations of core-localized TAEs in TFTR [93, 94, 95] and tornado modes in JT-60U [51]. Such modes were predicted to exist at low shear [96], consistent with the low shear value in this JET experiment. At high $P_{\text{ICRH}}$, the DBS beam reaches the cutoff surface at $\rho=0.251$ ($q\approx 0.92$) for $k_{\perp}\rho_{s}=3.78$. The power spectrum is broadband for fluctuations in the range $f\lesssim 600$ kHz, but also displays spectral peaks at the local Alfvén gap frequency $f_{\text{TAE}}\approx 290$ kHz and its multiples, which now extend to four (the difference in the background noise level of the curves in figure 3 originates from the difference in the absolute power launched from each of the corresponding DBS channels).
Figure 3: Power spectra $P_{s}(f)$ at low, medium, and high $P_{\text{ICRH}}$. At low $P_{\text{ICRH}}$, the spectrum is broadband, characteristic of drift-wave turbulence. At medium $P_{\text{ICRH}}$, note spectral peaks at ±$\pm f_{\text{TAE}}$ and ±$\pm 2f_{\text{TAE}}$. At high $P_{\text{ICRH}}$, $P_{s}(f)$ displays spectral peaks at $f_{\text{TAE}}$ and its multiples (up to four).
@@
Figure 4 shows the modulus of the bicoherence $|B_{s}(f_{1},f_{2})|$ for low, medium and high $P_{\text{ICRH}}$. At low $P_{\text{ICRH}}$ in 4.(a), the bicoherence is calculated for $255$ windows of period $T=0.1$ ms that are half-overlapping. Its modulus displays broadband phase-matched relationships between frequencies in the range $f_{1},f_{2}<100$ kHz, as expected from nonlinear interactions between turbulent eddies characteristic of drift-wave turbulence. The amplitude, however, remains low at $\approx 0.2$, and this suggests that nonlinear interactions between low-frequency drift-wave fluctuations are broadband and modest in the low-$P_{\text{ICRH}}$ phase. At medium $P_{\text{ICRH}}$ in 4.(b), the bicoherence is calculated for a total of $63$ half-overlapping windows of period $T=0.05$ ms. It displays broadband phase-matched relationships for multiples of the local Alfvén gap frequency $f_{\text{TAE}}$, the strongest of which are $(f_{1},f_{2})=\{(\pm f_{\text{TAE}},\pm f_{\text{TAE}})\}$ (amplitude $\approx 0.6$), which correspond to co-propagating TAEs that couple to generate a perturbation of frequency ±$\pm 2f_{\text{TAE}}$, as well as $(f_{1},f_{2})=\{(2f_{\text{TAE}},-f_{\text{TAE}}),(-f_{\text{TAE}},2f_{\text{TAE}})\}$, which correspond to perturbations of frequency $2f_{\text{TAE}}$ coupling to a counter-propagating perturbation at frequency $f_{\text{TAE}}$ that generates a co-propagating perturbation at $f_{\text{TAE}}$. The bicoherence also clearly exhibits phase-matched relationships between perturbations at multiples of $f_{\text{TAE}}$ and a low-frequency fluctuation, which we call the zero-frequency fluctuation in the rest of the manuscript. To our knowledge, this is the first experimental demonstration in a magnetically confined plasma that Alfvén modes living inside the fundamental toroidicity-induced frequency gap (and higher-frequency gaps) beat with counter-propagating perturbations to generate a zero-frequency fluctuation. The zero-frequency fluctuations observed here can naturally be interpreted as zonal modes driven by Alfvén eigenmodes, which have been predicted analytically [16, 17, 18, 19] and numerically [20, 22, 23, 24, 12, 11, 14]. Such zonal modes can suppress the turbulence driven by ITG, which would explain the increase of the ion temperature and its gradient (figure 2). At high $P_{\text{ICRH}}$ in 4.(c)-(d), the bicoherence is calculated for a total of $127$ half-overlapping windows of period $T=0.1$ ms. Similarly to 4.(b), the bicoherence displays broadband phase-matched relationships for multiples of $f_{\text{TAE}}$. Importantly, the highest-amplitude phase matching is now for the frequency pairs $(f_{1},f_{2})=\{(2f_{\text{TAE}},0),(2f_{\text{TAE}},-2f_{\text{TAE}}),(3f_{\text{TAE}},0),(3f_{\text{TAE}},-3f_{\text{TAE}})\}$ (amplitude $\approx 0.5$), i.e., the pairs that involve a zero-frequency fluctuation. This is different from the medium-$P_{\text{ICRH}}$ phase, in which the phase-matched relationships involving the zero-frequency fluctuation were subdominant to other mode-mode interactions. Interestingly, the modes at the fundamental Alfvén gap frequency $f_{\text{TAE}}$ display a weaker phase-matched relationship with the zero-frequency fluctuation. Higher multiples of the Alfvén frequency show weaker phase-matching, resulting in a peculiar grid structure.
Previous studies have observed bidirectional Alfvén modes in a tokamak [51, 57], but the nonlinear coupling with a zero-frequency zonal mode was not shown. TAEs and energetic-particle modes were also reported to interact nonlinearly with each other in [97], but they were not shown to couple nonlinearly to a zero-frequency zonal mode. In a recent publication [34], numerical simulations showed that TAEs could generate a strong zonal flow and stabilize the turbulence, leading to an enhanced-confinement DT plasma in JET. This is a promising result for extrapolation to burning plasmas, but it lacked the experimental verification of a zonal flow, which we provide here.
@@
Figure 4: Bicoherence spectra at low (a), medium (b), and high $P_{\text{ICRH}}$ (c)-(d). At low $P_{\text{ICRH}}$, we observe broadband phase-matched relationship in the range $f_{1},f_{2}<100$ kHz. At medium $P_{\text{ICRH}}$, the bicoherence displays phase-matched relationships for multiples of $f_{\text{TAE}}$ (vertical dashed lines), the strongest of which at ±$\pm f_{\text{TAE}}$ and ±$\pm 2f_{\text{TAE}}$. A weaker phase-matched relationship is observed between fluctuations at multiples of $f_{\text{TAE}}$ and a zero-frequency fluctuation. At high $P_{\text{ICRH}}$, the bicoherence displays phase-matched relationships for multiples of $f_{\text{TAE}}$. The strongest phase matching is for frequency pairs involving a zero-frequency fluctuation: $(f_{1},f_{2})=\{(\pm 2f_{\text{TAE}},0),(\pm 3f_{\text{TAE}},0)\}$. Panel (d) zooms into panel (c).
@@
Figures 4.(b)-(d) confirm that Alfvén modes living inside the toroidicity-induced Alfvén-frequency gap (and higher-frequency gaps) beat with counter-propagating Alfvén modes to generate a zero-frequency fluctuation that is consistent with a zonal mode. This behavior is correlated with an increase in the total energy-confinement factor ($H_{89,P}>1$), the main-ion temperature and its gradient (despite negligible direct ion heating), and a reduction in the main-ion heat diffusivity. This shows the strong beneficial effect that Alfvén eigenmodes can have on the confinement of the thermal plasma. This is to be constrasted with the stable-Alfvén-mode case (low $P_{\text{ICRH}}$), which exhibited low ion temperature and temperature gradient, and $H_{89,P}<1$. This is the first experimental confirmation in a magnetically confined plasma of a zero-frequency fluctuation that is pumped by Alfvén modes, and is correlated with hotter-than-expected main ions. It suggests that zonal modes driven by Alfvén modes can suppress the turbulence and lead to an overall improvement of energy confinement.
@@
Appendix A Linear gyrokinetic calculations
In this appendix, we confirm that the increase in the ion temperature and its gradient, the improvement of confinement, and the reduction in the ion heat diffusivity observed in the JET plasma presented in this letter cannot be explained by previous mechanisms of direct stabilization of turbulence by energetic particles [15]: dilution of the main plasma is negligible for the density of $n_{{}^{3}\text{He}}/n_{e}\approx 0.2-0.3\%$; stabilization by beta-related mechanisms is also negligible due to the low electron beta $\beta_{e}\lesssim 0.004$; and the fast-ion pressure and its gradient have a negligible effect on the equilibrium. The mechanisms of linear stabilization described in [98, 99] are not relevant either, since those effects are only significant at energetic-particle energies $E/T_{e}\approx 10$, in contrast to $E/T_{e}\gtrsim 10^{2}-10^{3}$ in the deep core of the plasma that we have studied. These linear, direct effects of the energetic particles have a negligible effect on the linear microinstability driven by the ion-temperature gradient (ITG), as we show in this appendix.
In order to prove that linear mechanisms are negligible in this JET configuration, we carry out linear micro-stability calculation of the most-unstable mode using the gyrokinetic code CGYRO [75] and the Pyrokinetics framework [76]. We extract the plasma parameters from the high-$P_{\text{ICRH}}$ case at the radial location $\rho\approx 0.3$, which overlaps with the experimental measurements of the zero-frequency fluctuations. Figure 5 shows the linear spectrum of the real frequency $\omega$ and growth rate $\gamma$ normalized by $c_{s}/a$, where $c_{s}=\sqrt{T_{e}/m_{D}}$ is the ion sound speed, $a$ is the minor radius, $m_{D}$ is the deuterium mass and $T_{e}$ is the local electron temperature. Two types of numerical calculations are performed: with a population of fast 3He ions, denoted “w/ EP” modelled by a Maxwellian of temperature equal to the mean energy of the fast particles, as provided by TRANSP-TORIC; and without the fast population of 3He, denoted “wo/ EP” (black). Both simulation types resolve thermal deuterium, hydrogen and electrons as gyrokinetic species. In the simulations “wo/ EP”, the value of the radial gradient of beta is scaled to maintain self-consistency of the species gradients with the magnetic equilibrium, and the density of the hydrogen is scaled to satisfy the quasi-neutrality condition (this modification is $<1\%$, due to the low density of 3He).
The linear calculation without fast 3He shows a dominant instability whose growth rate peaks at $k_{y}\rho_{s}\approx 0.6$, driven in the ion-diamagnetic drift direction $\omega>0$, consistent with the ITG mode. The simulations with fast 3He exhibit a very similar mode, but reduced by a factor of $<5\%$. This reduction is consistent with dilution, the effect of the gradient of beta on the equilibrium, and the linear kinetic effect discussed in [98, 99]. Additionally, at $k_{y}\rho_{s}\approx 0.01-0.05$ (see inset), a new destabilized mode appears, which is consistent with a TAE. Using the experimental plasma parameters, one can calculate the toroidal mode number $n$ that corresponds to the simulated $k_{y}\rho_{s}$. This leads to the prediction of unstable modes in the range of $n\approx 1-11$, consistent with the experimental observations of TAEs in figure 1 of the main manuscript (note that the simulation was run with the temperature gradient of 3He $a/L_{T_{{}^{3}\text{He}}}=15.7$ provided by TRANSP-TORIC, which is unrealistically large; $a/L_{T_{{}^{3}\text{He}}}$ is expected to be closer to marginal stability, which would reduce the growth rate and the range of toroidal mode numbers driven unstable). The real frequency of these modes is $\omega\approx 4c_{s}/a$. The dimensional value of the frequency using experimental parameters is $f\approx 260-290$ kHz, which is once more consistent with the experimental frequency range of the modes presented in figure 1 of the main manuscript.
Figure 5: Left panel: real-frequency $\omega$ vs. poloidal wavenumber $k_{y}\rho_{s}$ of the fastest-growing mode. Right panel: corresponding linear growth rate $\gamma$. Insets zoom in on the low-$k_{y}$ part of the spectrum. The case with a population of fast 3He ions is shown as magenta points, the case without as black ones.
The results of this appendix confirm that dilution, the effect of the gradient of beta on the equilibrium, and the linear kinetic effects of [98, 99] are negligible here: combined, they give rise to a $<5\%$ reduction of the linear growth rate of the fastest growing micro-instability. Having ruled out these possibilities leaves us with the effect of the zonal flow pumped by Alfvén modes as the only mechanism able to explain the ion-temperature increase and the improvement of confinement in this JET plasma. The experimental measurements of the zero-frequency fluctuation support this conclusion.